### v1

Abstract
Machine learning (ML) is a rapidly developing area of medicine that uses significant resources to apply computer science and statistics to medical issues. ML’s proponents laud its capacity to handle vast, complicated, and erratic medical data. It’s common knowledge that attackers might cause misclassification by deliberately creating inputs for machine learning classifiers. Research on adversarial examples has been extensively conducted in the field of computer vision applications. Healthcare systems are thought to be highly difficult because of the security and life-or-death considerations they include, and performance accuracy is very important. Recent arguments have suggested that adversarial attacks could be made against medical image analysis (MedIA) technologies because of the accompanying technology infrastructure and powerful financial incentives. Since the diagnosis will be the basis for important decisions, it is essential to assess how strong medical DNN tasks are against adversarial attacks. Simple adversarial attacks have been taken into account in several earlier studies. However, DNNs are susceptible to more risky and realistic attacks. The present paper covers recent proposed adversarial attack strategies against DNNs for medical imaging as well as countermeasures. In this study, we review current techniques for adversarial imaging attacks, detections. It also encompasses various facets of these techniques and offers suggestions for the robustness of neural networks to be improved in the future.
keywords:
Adversarial attack, Medical image, Deep Neural Network, Model safety, Robustness
1 Introduction
1.1 Adversarial Attack
Adversarial attacks encompass a set of techniques aimed at manipulating machine learning models by introducing well-crafted, often imperceptible alterations to input data. These modifications aim to deceive the model, leading to misclassifications or inaccurate predictions [1]. Adversarial attacks have the potential to compromise machine learning systems’ dependability and security, which is concerning for crucial applications such as cybersecurity, autonomous vehicles, and medical diagnosis. Of all the adversarial attack types, ”White-box attacks” and ”Black-box attacks” are the most common.
Ongoing research in machine learning security centers on the analysis of adversarial attacks and the development of defense strategies. Researchers continually explore novel attack tactics to identify potential weaknesses in AI systems while concurrently striving to construct robust models less susceptible to adversarial perturbations. Artificial intelligence (AI) modern technology is widely successful in fields like computer vision, natural language processing, and automated driving; however, its application in vital safety sectors is hindered by its susceptibility to adversarial attacks. Consequently, enhancing the resilience of AI systems against such attacks has become paramount for the progress of AI [2].
The rapid advancement of AI technologies has found diverse applications in numerous domains, from machine translation, speech recognition, and object identification to more intricate tasks like drug composition analysis [3, 4, 5, 6, 7]. Noteworthy applications include brain circuit construction [8], particle accelerator data analysis [9], [10], and DNA mutation impact analysis [11]. Since Szegedy et al.’s seminal work [12] highlighting neural networks’ vulnerability to adversarial attacks, research on adversarial technologies for artificial intelligence has steadily expanded. New techniques for adversarial attacks and strategies for mitigating them are consistently emerging.
The term ”adversarial attacks at the training stage” refers to actions taken by adversaries to manipulate the training dataset, input characteristics, or data labels during the training phase of the target model. This manipulation of the training dataset involves actions such as adding or deleting training data, as demonstrated by Barreno et al.’s approach [13]. During the testing phase, adversarial assaults are categorized into white-box attacks and black-box attacks [2]. In white-box situations, attackers have access to the parameters, techniques, and structure of the target model, enabling them to craft adversarial samples based on this information.
1.2 The Adversary’s Objective: Evasion attack versus poisoning assault
”Poisoning attacks” are assault methods that allow an attacker to add or change a large number of fake samples to the training set of a DNN algorithm. The trained classifier may perform poorly as a result of these bogus data. They may have poor accuracy [14] conversely make imprecise predictions on a few samples for analysis [15]. The classifiers used in evasion attacks are fixed and often work well on safe testing samples. The adversaries are unable to reform the classifier’s settings or parameters, but they do generate some fictitious instances that the classifier is unable to distinguish.
1.3 Deep Neural Networks: Adversarial Attacks
Regarding medical image processing use cases such as diagnosing cancer and lesion identification, deep neural networks, also known as DNNs, are rapidly gaining popularity [16]. Nonetheless, a new study suggests that adversarial examples/attacks with tiny, barely noticeable changes can weaken medical deep learning systems. There are presently safety concerns associated with the usage of these devices in clinical settings [16]. Deep neural networks (DNN) are increasingly well-liked and effective at a variety of machine learning requisitions. They have been employed with surprising effectiveness in a variety of recognition issues in the fields of pictures, graphs, text, and voice [1]. They are able to identify things in images with accuracy that is almost human [17], [18]. Additionally, they are employed in speech recognition [19], natural language processing [20], and gaming [21].
Figure 1: The attack of Biggio’s SVM classifier for letter recognition. [1]
Regarding the data set provided by MNIST, Biggio et al.[22] produce adversarial instances first, focusing on traditional machine learning classifiers like SVMs and 3-layer neural networks with full connectivity which is shown in Figure 1. In order to trick the classifier, it optimizes the discriminant function. As an illustration, consider a linear SVM classifier on the MNIST dataset.
1.4 Examining examples of opposition in the real world
The research [23] put decals to traffic signs that pose a serious threat to autonomous vehicles’ sign recognition technology. These hostile objects are especially detrimental to deep learning systems since they can interfere directly with numerous real-world DNN usages, such facial recognition and autonomous cars. By assessing the adversarial images (FGSM, BIM) produced to see if they are ”robust” against changes in natural circumstances (e.g., shifting viewpoint, illumination, etc.), the authors of the work [24] investigate the viability of creating tangible adversarial objects. Robust in this context means that even after transformation, the produced pictures are still antagonistic. The experiment’s findings show that many of these adversarial examples—particularly those produced by FGSM—remain antagonistic to the classifier even after transformation. The findings imply that antagonistic real-world objects could trick the sensor in many situations.
1.5 Medical Image Under Artificial Intelligence Attack
Researchers have access to strong models of developing science and technology thanks to deep learning. Convolutional neural networks, also known as CNNs, are among the most significant categories of deep learning (DL) frameworks for the analysis and processing of images because of their exceptional ability to learn valuable information. The processing of the human organism utilizing different picture modalities for therapeutic, diagnostic, and health surveillance purposes is known as medical image analysis [25]. The advancement of computer vision through the use of deep neural networks addresses issues that were not well-solved by traditional image processing methods. Beinfeld et al. [26] asserted that \$385 spent on medical imaging results in a savings of almost \$3000. MRI, CT scans, ultrasound (US), and X-rays are the most frequently used image modalities. Figure 2 shows how adversarial assaults may be used to arbitrary alter diagnosis outcomes in three different medical picture datasets: Dermoscopy [27], X-ray of the chest [28], and Fundoscopy [29].
Figure 2: Instances of adversarial approaches designed by the PGD (Projected Gradient Descent) to deceive DNNs developed using datasets of medical images include dermoscopy [27] (third row), chest x-ray [28], and fundoscopy [29] (first row, DR=diabetic retinopathy). Normal images on the left, adversarial perturbations in the middle, and adversarial images on the right. The left-bottom tag indicates the projected class, and green or red indicates whether the predictions were accurate or not [16].
Hu et al. [1] assert that in order to develop more reliable models, it is crucial to investigate the reasons why adversarial cases emerge and to comprehend deep learning models better. Depending on the knowledge of the enemy, attacks can be categorized into three kinds. The saliency (or attention) map of a picture that was input shows the regions that significantly change the model’s output based on the gradients of the categorization loss with regard to the input [16]. We can see that some medical photos have highly concentrated regions that are noticeably larger. This could indicate that the rich biological textures included in medical photos occasionally entice the DNN model to focus more on aspects unrelated to the diagnosis. Tiny adjustments in these high-focus areas can significantly affect the model’s output.
Due to several factors, including High-Dimensional Data, High-Dimensional Data, Transferability of Adversarial Examples, Complex Decision Boundaries, Black-Box Attacks, Safety-Critical Applications, and Safety-Critical Applications, Medical Image Deep Neural Network (DNN) models can be relatively easier to attack when compared to some other domains. Despite these difficulties, experts in the field are actively attempting to create more reliable and safe DNN models for medical images. To make these models more resistant to hostile attacks, strategies like adversarial training, input preprocessing, and defensive distillation are being investigated. To further enhance the security and dependability of AI systems in healthcare applications, consistent evaluation frameworks for adversarial robustness in medical imaging must be developed.
Byra et al. [30] proposed an attack strategy on ultrasonography (US) imaging for liver fatty tissue. Radio-frequency signals are used to rebuild US pictures, and the reconstruction technique was subjected to a zeroth-order optimization attack [31]. The InceptionResNetV2 model was used in the studies, and the assault resulted in a 48% loss in model accuracy. Ozbulak et al. [32] proposed an attack specifically designed for medical picture segmentation called the adaptive segmentation masking attack (ASMA). The suggested attack offers significant intersection-over-union (IoU) degradation and produces nearly undetectable samples for most portions. Because the U-Net framework is among the most well-known models for clinical picture segmentation, they employed it in the trials. The datasets used were for segmentation of glaucoma optic disk [33] and ISIC skin lesion segmentation [34]. Chen et al. published a method for fabricating hostile cases to thwart medical picture segmentation [35]. In order to simulate anatomical and intensity fluctuations, geometrical deformations are used to create the adversarial examples. By attempting to partition organs from abdominal CT scans using a U-Net model, they tested the effectiveness of these examples. With reference to the Dice score measure, they successfully reduced it significantly across all organs. The kidneys and pancreas, however, require a higher level of disturbance and are more arduous to assault than the liver as well as spleen.
Figure 3: Clean image, bias field noise, and diagnosis following application of bias field noise are shown in (a), (b), and (c), respectively [36].
Tian et al. [36] examined the phenomena of bias field, which can arise from incorrectly capturing a medical image and jeopardize a DNN’s efficacy, as shown in Figure 3. Motivated by adversarial attacks, the authors created an adversarial-smooth bias discipline approach to trick a model. The ResNet50, MobileNet, among others and DenseNet121 models were employed to fine-tune the chest X-ray dataset that was used in their study. They looked at this attack’s transferability and white-box attacks. Comparing the suggested attack to other cutting-edge white-box attacks, it exhibited a greater attack accuracy on transferability.
This review’s major organization is as follows: We offer some significant adversarial notions on medical images in Section 2. Additionally, it provides a detailed overview of earlier research. We explore a few strategies in Section 3 with regard to the picture classification scenario. We utilize Section 4 to quickly summarize some findings from earlier research that tries to explain the phenomena of antagonistic examples on medical data. The overview is concluded in Section 5.
2 Background Literature Review
Adversarial instances are intentionally created data inputs that are intended to degrade a machine learning model’s performance. When researchers examined strategies used by spammers to evade spam filters in 2004, they unofficially coined the term ”adversarial inputs” [37]. These misleading examples are usually produced by deliberately faking real data, such spam advertising messages, in order to trick the computer system that analyzes it. Alterations can be applied to text data, like spam, by introducing innocuous text or modifying frequently used phrases in malicious communications with synonyms. To fortify algorithms against adversarial attacks, researchers have explored various strategies, including training algorithms on adversarial samples and employing sophisticated data processing techniques to minimize the susceptibility to manipulation. The quest for fully robust models in machine learning aims to accelerate the development of algorithms capable of making decisions based on consistent explanations, with promising early efforts in this direction [38].
In the realm of healthcare, medical claims codes play a crucial role in determining the amount spent on a patient visit post-payer approval. Payers often assess these claims using automated fraud detectors increasingly driven by machine learning. Historically, healthcare providers have shaped payer records of patient visits, including the associated codes, to influence the algorithmic outputs of payers [38]. Medical fraud, a market valued at \$250 billion, exemplifies the extreme end of this strategic tailoring of patient presentations. While some practitioners may overtly fabricate medical claims, patient data falsification often takes more covert forms. For instance, consistently submitting codes for billing services identical to, but more expensive than, those actually provided is known as intentional upcoding. Here, Table 1 shows past exclusive past research works where authors did some fantastic work for medical images to defend adversarial attacks. Different researcher proposed different frameworks to overcome this problem and to suggest a new thought for revolution.
2.1 Elements of a Cyber Adversary’s Assault
Figure 4: A showcase of executing adversarial attacks on various medical AI systems without resorting to overtly deceptive data manipulation [39].
In some instances, subtle adjustments to billing codes blur the boundary between fraudulent activities and conscientious best practices. One notable illustration may be found in the guidelines on the Endocrine Society website, which advise medical professionals not to bill for metabolic syndrome, or International Classification of Diseases (ICD) code 277.77, in situations of obesity. Coverage denial is likely to occur with this particular code/condition combo in Figure 4. The Society recommends an alternative approach, suggesting billing for the individual codes related to specific disorders comprising the metabolic syndrome, such as hypertension.
Table 1: Existing adversarial attack research on medical Image
Reference Year Framework Contribution
[40] 2021 Propose a framework to defend against adversarial training, speckle-noise attacks, and maintain accurate classification labels in recognizing diabetic retinopathy through analyzing retinal fundus images. a) accuracy of 99%; b) defensive model’s robustness ; c) a system that includes a fresh SN attack.
[41] 2022 Explain the use of gradient-free trained sign activation systems in medical imaging AI systems to identify and thwart hostile attacks. The model performs better and even twice as well as the finest.  a) Owing to significant distortion and excellent transferability accuracy, the model’s average success rate in adversarial case classification is 88.89%; b) This model defends against adversarial attacks.
[42] 2020 This method detects diverse adversarial attacks while preserving user anonymity and classification performance, making it a versatile addition for deep learning-based medical imaging systems to enhance resilience. a) The chest X-ray dataset exhibits robust performance across diverse configurations, contributing to the heightened safety of deep learning-based medical image classification systems; b) To improve medical image categorization systems based on deep learning security.
[43] 2022 In weakly-supervised clinical tasks, CNNs and ViTs exhibit vulnerability to both white- and black-box adversarial attacks with comparable baseline performance. a)Compared to CNNs, aViTs have a larger latent portrayal of clinically meaningful categories; b) Findings align with prior theoretical investigations and provide tangible evidence of ViTs’ capacity to grasp computational pathology; c) This suggests that computational pathology AI models will be deployed widely.
[…]
2.2 The components of an adversarial attack
Figure 5: Distribution of chosen articles on a yearly basis Pie chart; line graph [44].
Presenting a thorough overview of the various adversarial attack strategies and defense techniques is the aim of the study [44]. We first discuss the theoretical underpinnings, practices, and practical uses of adversarial attack strategies. Then, a brief discussion of the research on defense strategies obscuring the boundary of the large field follows. A few selected articles are published year-by-year from 2012 to 2021, as seen in Figure 5. This work aims to provide thorough taxonomies, evaluations of harmful assaults, and protections against the full DL pipeline. In this context, the numerous attack and defense tactics that have been developed over the previous two years have been categorized, with a focus on the clinical deep learning systems that are susceptible to adversarial attacks.
Several defensive strategies have been put forth in an effort to neutralize possible threats. A popular method in natural imaging called adversarial training adds adversarial images to the training dataset in order to make Convolutional neural network, or CNN, models more resilient. Nevertheless, this method is not the best for datasets pertaining to medical imaging, since the addition of different adversarial images could considerably reduce classification accuracy. A robust detection technique for malicious images is presented in the work of Li et al. [42], successfully defending against attacks on deep learning-driven medical image labeling systems. In a different study [43], scientists compared CNN performance with Vision Transformers (ViTs) to determine how resilient CNNs are to different types of attacks in computational pathology. The authors developed robust neural network models, evaluating their efficacy against both white- and black-box attacks. The structures of attacks for both models were scrutinized, with an exploration of the underlying factors influencing their performance. The study’s findings were validated through two clinically relevant classification tasks involving distinct patient groups. Preceding Ma et al.’s research [16], ambiguity surrounded medical image adversarial attacks (AAs), limiting adversarial machine learning analyses to natural images. Unlike natural images, medical images may contain domain-specific elements, impacting medical deep-learning systems susceptible to AAs. AAs possess the potential to manipulate diagnoses and outcomes, emphasizing the critical need for a robust healthcare infrastructure to mitigate potential risks from malicious attacks.
3 Materials and Methods
3.1 Medical Adversarial Defensive Strategies
Various defense models, including techniques such as input denoising, input gradients regularization, and adversarial training, have been devised. However, recent attacks often manage to evade or partially circumvent these protective measures. This paper undertakes a comprehensive exploration of adversarial attack and defense methods within the domain of medical image analysis. It introduces a novel taxonomy based on the application context for a family of techniques, as outlined in Dong et al.’s work [45]. The author establishes a unified theoretical framework encompassing diverse adversarial attack and defense strategies tailored for medical imaging applications.
The primary emphasis in most research efforts on adversarial attacks in medical image processing is on the white-box scenario. These studies, characterized by a comprehensive understanding of medical Deep Neural Networks (DNNs), concentrate on the susceptibility of computer-aided diagnosis models across various medical imaging applications. In the execution of adversarial attacks, attackers may utilize the target diagnosis DNN as a locally deployed model. In addition to white-box adversarial attacks for medical classification tasks, scholars have delved into vulnerabilities associated with various medical imaging tasks. Notably, these publications focus predominantly on medical segmentation, as evident in works such as Chen et al.’s [35] and Ozbulak et al.’s [32].
For natural photos, semi-white-box (Gray-box) attacks have been extensively researched [46]. However, there are only a few academic publications [47], [48] that address this attack scenario for medical image processing. The semi-white-box adversarial approach typically consists of two stages: 1) To create adversarial instances against the target DNN model, the attacker trains a generative model. The attacker has complete access to the target model during training, including backward propagation gradients. 2) Instead of needing to know anything about the target model, as would be the case in a completely black-box scenario, the adversary generator can directly obtain adversarial examples against the target model with the input of authentic photos during the application stage.
Presently utilized white-box adversarial attacks often require multiple backward gradients of the target model. In simpler terms, attackers generate comparable adversarial samples by treating the target Deep Neural Network (DNN) as if it were a locally deployed model. However, due to its reliance on an in-depth understanding of the DNN model to execute an attack, the white-box approach may be unreliable in real-world scenarios. Conversely, the general black-box scenario may offer a more suitable environment for simulating real-world adversarial attacks, with numerous proposals focusing on investigating black-box assaults for natural images.
The outlined solutions usually involve multiple queries to the black-box model or depend on having a comprehensive understanding of the desired diagnosis model. However, in many real-world situations, attackers may lack direct access to the target diagnosis model. The restricted black-box (no-box) configuration, specifically representing the most challenging scenario for real-world adversarial assaults, may be more covertly dangerous even without querying the target black-box DNN. The transferability [49] of adversarial instances across different DNN models fundamentally determines the success of no-box attacks. For example, a no-box attacker can generate adversarial images based on a locally deployed surrogate model, transferring them directly to target medical diagnosis systems. Restricted black-box adversarial attacks are deemed more subtly perilous for natural vision tasks according to a renowned study.
Considering the substantial risks to the healthcare sector, numerous defense strategies have been proposed to counter medical adversarial attacks. In order to overcome the weaknesses shown by adversarial examples, a number of methods have been developed to provide trustworthy deep learning-based systems for natural pictures [49]. The development of accurate computer-aided diagnosis models for clinical applications not only contributes to trustworthy healthcare services for millions but also underscores the importance of investigating adversarially robust models within the realm of medical image analysis.
In their work [45], the authors succinctly summarize studies on adversarial defense in the context of medical image analysis. Various attack methods in the realm of medical images serve as robustness evaluation criteria for adversarial defense, complementing the extensive range of adversarial attack methods mentioned. The minimal separation between adversarial examples and the model’s decision boundary signifies the model’s resistance to disturbances. The goal of adversarial training is to mitigate subsequent losses.
Minimize the
$L(f(x_{\text{adv}},y)+\lambda\cdot L(x,x_{\text{adv}}))$ (1)
where $L(f(x_{\text{adv}},y)$ is the classification loss, $L(x,x_{\text{adv}})$ is the perturbation size, and $\lambda$ is a trade-off hyperparameter.
Transforming the input data into an alternative feature space that diminishes the impact of adversarial perturbations on the model. A technique aimed at enhancing resilience against adversarial attacks involves introducing noise to the features or intermediate representations that characterize the input data.
3.2 Adversarial Training
Most medical adversarial defense techniques focus on using adversarial training to create reliable diagnoses systems. A significant fraction of the works among them go beyond current techniques for natural image adversarial training to tasks relating to medical classification [50], [51], Vatian et al. [52] looked into opposing examples for medical imaging and tried a number of defense strategies to oppose these nefarious representations.
3.3 Adversarial Detection
Adversary detection seeks to identify adversary cases from input examples during the application stage, as opposed to developing robustness during the training stage of computer-aided diagnosis models.Many adversarial detection techniques have been put forth in the field of medical image analysis to stop additional misdiagnosis caused by adversarial samples [42]. In particular, it is possible to think about medical adversarial detection as an anomaly detection issue that can be resolved by combining explainability approaches [53].
3.4 Image-level Pre-processing
A clean image and the related adversarial perturbation make up an adversarial image in general. Meanwhile, it has been shown that DNNs can perform well on clean images while still being vulnerable to adversarial examples [12]. Denoising the adversarial example to remove the perturbation component can therefore help make the subsequent network diagnostic easier. Image-level pre-processing can be useful and secure in the context of biomedical image analysis because it does not require re-training or modifying medical models.
3.5 Improvement of Features
The difference in robustness between human and machine vision is attributed to adversarial examples linked to non-robust features extracted from specific patterns in the data distribution [54]. Consequently, enhancing feature representation is crucial for the development of robust inference systems. In this investigation, we characterize feature augmentation as the modification of architectures or mapping functions. Various techniques aimed at improving features have significantly enhanced the robustness of medical classification models [55].
3.6 Distillation of Knowledge
In the field of machine learning, knowledge distillation is a helpful technique for transferring learned information from a complex (teacher) model to a simple (student) model. Therefore, the situation where the network structures for the teacher and learners are identical is specifically referred to as self-distillation. Furthermore, a great deal of research has been done on the natural imaging domain using adversarial knowledge distillation, which moves the adversarial resilience from a heavy teacher model to a view student model [56].
4 Result and Discussion
This study utilizes four publicly accessible benchmark datasets, as detailed in [45], to investigate adversarial attack and defense in the context of medical image processing (illustrated in Figure 6). Firstly, the Messidor1 dataset comprises 1,200 eye fundus color images for detecting diabetic retinopathy across four classes based on retinopathy grade. Secondly, the International Skin Imaging Collaboration (ISIC) dataset consists of 2,750 dermoscopic images categorized into three classes for skin lesion classification and segmentation. Thirdly, the ChestX-ray 14 dataset comprises 112,120 frontal-view X-ray images representing 14 thorax diseases. Lastly, the COVID-19 database incorporates 21,165 chest X-ray images accompanied by segmentation-capable lung masks.
Figure 6: Medical adversarial examples with predictions for a range of perturbation sizes. For visualization, the created segmentation masks are placed on top of the source photos [45].
The study conducted by the authors [57] employs experiments to demonstrate the following: 1) The SSAT module significantly enhances the adversarial robustness of the model without compromising the classification accuracy of clean images. 2) The UAD module effectively identifies and excludes a majority of successful adversarial examples. 3) In comparison to other existing AI systems, their medical imaging AI solution (UAD + SSAT) minimizes the risk of adversarial attacks. The research utilizes a publicly available dataset of retinal OCT images, employing the ”white-box” setting as the most challenging threat for assessing class prediction performance. Across all white-box attack conditions, the authors establish that SSAT consistently outperforms alternative baselines while maintaining comparable or superior performance for clean image classification. Furthermore, the authors demonstrate that the combined approach of UAD complementing SSAT results in the lowest adversarial risk according to the new measure presented. Regardless of the training techniques employed, systems based on UAD consistently exhibit reduced risks compared to those that do not incorporate this defense mechanism.
Figure 7: An outline of our study’s methodology. [57]
An AI-CAD model depicted in Figure 7 was initially trained to manipulate diagnosis-sensitive elements within images, such as adding or removing malignant tissue. The model’s effectiveness was then assessed using adversarial images produced by a GAN model. To determine how well human specialists could visually distinguish images produced by the GAN, a reader study was carried out.
In a study by Zhou et al. [58], the responses of an AI-CAD model to adversarial attacks on GAN-generated mammography images were investigated. This involved introducing malignant tissue into healthy images and removing cancerous tissue from images affected by cancer. The study also evaluated the proficiency of experienced radiologists in visually identifying such adversarial images, both before and after instructional intervention. The study cohort, provided by the University of Pittsburgh Medical Center, included 1284 women, and 4346 mammography images were collected. Among these, 366 patients had biopsy-proven malignant breast cancer, while 918 patients had breast cancer assessed as benign (including benign signs).
Following training, the researchers assessed the model’s classification accuracy concerning both the original genuine test data and its equivalent GAN-generated adversarial counterparts. Two GAN-trained U-Net23 models were employed to generate adversarial images, flipping labels in the test set. The high-resolution (1728 x 1408) photographs indicated the categorization impact of the AI-CAD model on the test data. The AI-CAD model achieved an AUC of 0.82 on the test set, consisting of 364 genuine negative samples and 74 real positive samples. The adversarial GAN-generated images (with flipped labels) in the test set yielded an AUC of 0.94. These AUC values suggested that the adversarial collection of images effectively deceived the AI-CAD model. Additionally, 59.5% (44 out of 74 cases) of genuine positive images with a classification accuracy threshold of 0.5 were correctly labeled as positive, whereas 95.5% (42 out of 44 cases) of adversarial samples created by GANs successfully tricked the system by being classified as negative.
Figure 8: The Mean Structural Similarity Index Measure (SSIM) was computed between the authentic images in the test sets and the adversarial instances generated using FGSM or PGD at different perturbation levels. The SSIM values displayed represent the average across two model architectures (Inception-v3 and Densenet-121) [59].
Figure 8 displays the mean SSIM values for FGSM and PGD assaults across all photos. The Supplementary Material includes the SSIM data for every model. As can be shown, the effects of the identical disturbance applied to various imaging modalities on human visual perceptibility with the observed SSIM varies. The radiology images most clearly showed adversarial disturbances, with = 0.02 producing an already apparent, but very modest perturbation. At the same perturbation level for the ophthalmology and pathology photos, perturbations were nearly undetectable and became apparent with larger epsilon values. The authors examine black-box adversarial attacks on deep learning in healthcare imaging. [59]. They investigate three medical imaging sectors where deep learning algorithms are vulnerable. In all three datasets, perturbations calculated by FGSM showed lower SSIM than those calculated using PGD. This is a predicted outcome given that PGD optimizes perturbations based on their size and influence on model predictions.We opted to document attacks in our subsequent analyses employing 0.02, as this represented the maximum perturbation level that remained effective across all applications and attack strategies. Additionally, it exhibited greater transferability than an epsilon value of 0.01 in the majority of the examined applications.
Figure 9: In a range of black-box settings, such as target and surrogate pre-trained on ImageNet, target and surrogate at random initialized, target and replacement randomly initialized plus surrogate developed using a different and reduced dataset (d2/2), FGSM (=0.02) was used to create the original images, adversarial images, and corresponding adversarial noise. The aforementioned shows the average receiver operating characteristic curve (AUC) for both clean and black-box scenarios for each configuration. Diabetic retinopathy (DR) is misclassified in a red frame and correctly classified as referable or non-referable in a green frame. The adversarial noise  represents the difference between the initial and adversarial image.citeadversarial.bortsova2021.
Figure 9 shows examples taken from the ophthalmology collection that demonstrate the transferability of assaults when the target and surrogate are randomly started and when both are pre-trained on ImageNet. Whether the target and surrogate models had the same architecture or a different one did not affect the findings.
5 Conclusion
Diagnostic imaging AI systems based on computer vision are increasingly being used as models for classifying and segmenting diseases. The scaled-up use of medical imaging AI systems has given rise to serious safety issues because to DNNs’ susceptibility to adversarial samples. Recently, a number of ways have been put out to increase the efficacy of medical image defense tactics. Although numerous protection strategies have been put forth, there are still reservations over the use of medical deep learning techniques. This outcome arises from certain constraints within medical imaging, including a scarcity of high-quality image datasets and labeled data in comparison to more abundant datasets available for natural images. The efficacy of adversarial defenses may not be universally applicable in this context. The findings underscore the importance of exercising caution in the design of Deep Neural Networks (DNNs) for medical imaging and their subsequent real-world applications.
Declarations
Funding
A portion of the funding for this research comes from DOEd Grant P116Z220008 (1) and NSF Grant No. 2306109. The author(s) expresses all opinions, findings, and conclusions; the sponsor(s) does not necessarily agree with them.
Conflict of interest
The authors do not declare any conflicts of interest.
Ethics approval
There is no original research involving human subjects, animal subjects, or sensitive data in this review publication.
Consent to participate
The studies and data included in this review were sourced from publicly available and previously published literature.
Consent for publication
All authors listed on this review paper have reviewed and approved the final manuscript for submission to Machine Learning. Each author has contributed significantly to the research, writing, and revision of the paper. All authors have read and agreed to the content presented in this review, and their consent for publication is hereby provided.
Availability of data and materials
The data and materials used can be found in the references of this work.
Code availability
The review is based on a comprehensive analysis of publicly available literature, and any references to specific methodologies or algorithms are attributed to the respective original sources. We have strived to provide clear citations and references to enable readers to locate the original works for further examination.
Authors’ contributions
Conceptualization: A.B., M.A.A.N.; Methodology: A.B., M.A.A.N.; Formal analysis and investigation: A.B., M.A.A.N.; Writing - original draft preparation: A.B., M.A.A.N., R.G., A.R.; Writing - review and editing: M.A.A.N., A.R.; Funding acquisition: K.D.G; Resources: R.G.; Supervision: K.D.G., R.G. All authors have read and agreed to the published version of the manuscript.